# Partial Symmetry Enforced Attention Decomposition (PSEAD): A Group-Theoretic Framework for Equivariant Transformers in Biological Systems

### v1

@@ Abstract
Biological systems, from the intricacies of genomic sequences and protein structures to the dynamic processes of cellular signaling and morphogenesis, are fundamentally characterized by inherent symmetries. These symmetries, often local, partial, or approximate, manifest as repeating patterns, conserved motifs, or shared structural principles. Traditional machine learning paradigms frequently falter in effectively capturing or exploiting these deep-seated inductive biases, either by ignoring them entirely or by imposing restrictive global equivariance constraints that fail to mirror the nuanced reality of biological phenomena.
This research introduces the Theory of Partial Symmetry Enforced Attention Decomposition (PSEAD), a novel, rigorous group-theoretic framework designed to seamlessly integrate local symmetry awareness into the core architecture of self-attention mechanisms within Transformer models. We formalize the concept of local permutation subgroup actions on windows of biological data, proving that under such actions, the attention mechanism naturally decomposes into a direct sum of orthogonal irreducible components. Critically, these components are intrinsically aligned with the irreducible representations of the acting permutation subgroup, thereby providing a powerful mathematical basis for disentangling symmetric and asymmetric features.
We demonstrate the analytical derivation and computational implementation of PSEAD, illustrating its capacity to imbue Transformer architectures with a profound understanding of localized symmetries. Through a multi-faceted approach encompassing theoretical proofs, algorithmic designs, and preliminary empirical validations, we show that PSEAD offers substantial advantages. These include enhanced generalization capabilities to novel biological motifs exhibiting similar partial symmetries, unprecedented interpretability by allowing direct visualization and analysis of attention contributions from different symmetry channels, and significant computational efficiency gains by focusing representational capacity on relevant symmetric subspaces. Beyond static data analysis, we extend PSEAD’s applicability to dynamic biological processes within reinforcement learning paradigms, showcasing its potential to accelerate the discovery and optimization of biologically meaningful policies in complex environments like protein folding and drug discovery. This work lays the groundwork for a new generation of biologically informed, symmetry-aware artificial intelligence models.
@@ 2.1 Group Actions, Representations, and Equivariance
[…]
@@ 3.2 Case Study: Protein Folding as a Partially Symmetric Markov Decision Process (MDP)
By leveraging the symmetry inductive bias, the agent should require significantly fewer environmental interactions (e.g., protein folding simulations) to converge to an optimal or near-optimal folding policy. This is critical for computationally expensive biological simulations.
The irrep-wise attention mechanisms within PSEAD can highlight which specific local symmetries are most salient for particular folding steps or for achieving stable intermediate structures. For instance, attention heads corresponding to reflectional irreps might be highly active when forming $\beta$-sheets, while rotational irreps might be crucial for helical packing. This offers mechanistic interpretability of the folding process.
@@
# Please click this link to view:
https://github.com/DanielAyomide-git/psead/blob/main/listings/listing1.py
Listing 1: Simplified PSEAD Policy Network for Protein Folding
Information on how to install PSEAD python module or future code updates can be found here:
https://github.com/DanielAyomide-git/psead
@@ 4.1 Case Study 1: DNA Palindromes and Regulatory Motifs with $\mathbb{Z}_{2}$ Symmetry
We implement the PSEADAttention module with group=CyclicGroup(2).
import torch
import torch.nn as nn
from psead.attention import PSEADAttention
from psead.groups import CyclicGroup
# Example: Encode a DNA window (e.g., 10-mer) into numerical features
def encode_dna_window(dna_sequence: str, char_to_vec_map: dict):
# Example: one-hot encoding or learned embeddings
# ’A’: [1,0,0,0], ’C’: [0,1,0,0], ...
encoded_seq = torch.tensor(
[char_to_vec_map[base] for base in dna_sequence], dtype=torch.float32
return encoded_seq.unsqueeze(0) # Add batch dimension
# Assume a simple character to vector map
char_to_vec = {
"A": [1.0, 0, 0, 0],
"C": [0, 1.0, 0, 0],
"G": [0, 0, 1.0, 0],
"T": [0, 0, 0, 1.0],
feature_dim = 4 # One-hot encoding dimension
# Example DNA window with Z2 symmetry (palindrome)
dna_window_seq_2mer = "AG" # k=2 for Z2 example
dna_window_features = encode_dna_window(
dna_window_seq_2mer, char_to_vec
) # shape (1, 2, 4)
# Instantiate PSEADAttention for Z2 group
psead_attn_layer = PSEADAttention(group=CyclicGroup(2), input_dim=feature_dim)
# Compute the PSEAD-augmented attention output
out_psead_attn = psead_attn_layer(
dna_window_features
) # Corrected: Pass only one input tensor
print(f"Input shape: {dna_window_features.shape}")
print(f"Output shape from PSEAD attention: {out_psead_attn.shape}")
Listing 2: PSEAD for DNA Palindrome Analysis
The module computes attention for each irrep subspace.
@@ 4.2 Case Study 2: Protein Motifs with Dihedral ($D_{n}$) or Cyclic ($C_{n}$) Symmetries
Implementation: The PSEADAttention module would be initialized with group=DihedralGroup(n) or CyclicGroup(n). The window_size would correspond to the number of residues or structural elements within the symmetric motif.
import torch
import torch.nn as nn
from psead.attention import PSEADAttention
from psead.groups import DihedralGroup, CyclicGroup
# Example: Protein residue features for a D2-symmetric motif (k=2)
# For a general D_n, k would be n. This example uses D2 with k=2.
window_size_protein = 2 # Number of ’elements’ in our D2-symmetric motif
feature_dim_protein = 128 # Feature dimension for each element
protein_features = torch.randn(1, window_size_protein, feature_dim_protein) # (batch_size, k, feature_dim)
# Define Q, K, V (simplified)
Q_p = K_p = V_p = protein_features
# Instantiate PSEADAttention for Dihedral Group D2
psead_attn_d2 = PSEADAttention(group=DihedralGroup(2), input_dim=feature_dim_protein)
# Compute the PSEAD-augmented attention output
out_psead_attn_d2 = psead_attn_d2(protein_features) # Corrected: Pass only one input tensor
print(f"Input protein feature shape: {protein_features.shape}")
print(f"Output protein feature shape from PSEAD attention (D2): {out_psead_attn_d2.shape}")
Listing 3: PSEAD for Protein Motif Analysis (D2 Example)

### v2

@@ Abstract
Biological systems—from genomic sequences and protein structures to cellular signaling and morphogenesis—are shaped by local, partial, and approximate symmetries such as repeating patterns and conserved motifs. Traditional machine learning models often fail to capture these inductive biases, either ignoring them or enforcing overly restrictive global equivariance.
We introduce the Theory of Partial Symmetry Enforced Attention Decomposition (PSEAD), a group-theoretic framework that embeds local symmetry awareness into Transformer self-attention. By formalizing local permutation subgroup actions on data windows, we prove that attention decomposes into orthogonal irreducible components aligned with the subgroup’s representations. This provides a principled method for separating symmetric and asymmetric features.
We present the theory and algorithms demonstrating that PSEAD improves generalization to novel motifs with similar partial symmetries, enhances interpretability through symmetry-channel analysis, and increases computational efficiency by focusing on relevant symmetric subspaces. We further extend the framework to dynamic biological processes in reinforcement learning, highlighting applications to protein folding and drug discovery. PSEAD establishes a foundation for biologically informed, symmetry-aware AI models. Source code available here: https://github.com/DanielAyomide-git/psead
@@ 2.1 Group Actions, Representations, and Equivariance
Since this holds for all $i$, we have $\rho(h_{1})\rho(h_{2})(x)=\rho(h_{1}h_{2})(x)$. Therefore, $\rho(h_{1})\rho(h_{2})=\rho(h_{1}h_{2})$, confirming that $\rho$ is a group homomorphism. $\quad\blacksquare$
@@ 3.2 Case Study: Protein Folding as a Partially Symmetric Markov Decision Process (MDP)
By leveraging the symmetry inductive bias, the agent should require significantly fewer environmental interactions (e.g., protein folding simulations) to converge to an optimal or near-optimal folding policy. This is critical for computationally expensive biological simulations. The irrep-wise attention mechanisms within PSEAD can highlight which specific local symmetries are most salient for particular folding steps or for achieving stable intermediate structures. For instance, attention heads corresponding to reflectional irreps might be highly active when forming $\beta$-sheets, while rotational irreps might be crucial for helical packing. This offers mechanistic interpretability of the folding process.